2.) SPIs preferred stock pays a $1.00 dividend per quarter. The required return, r, is 2.5% per quarter. a) First assume the dividend will not grow. What is the price of a share of SPI preferred stock? b) What is the price if the dividend were expected to grow at 0.5% per quarter?
Answer #1

2)

a)Price = Dividend /required return

             = 1 /.025

               = $40 per share

b) price= D0(1+g)/(Rs-g)

           = 1(1+.005)/(.025-.005)

              =1 *1.005 /.02

               =50.25 per share